# Binary Black Hole Waveforms from High-Resolution GR-Athena++ Simulations

### v1

@@ Appendix F Time domain comparisons with SXS
In this appendix we report time domain comparisons similar to the one presented in the main text (Fig. 6) for all mass ratios considered in this work. The results are shown in Figs. 19, 20 and 21. Quantitative figures such as amplitude and phase differences are comparable with the ones discussed in the main portion of the text for the $q=2$ system. Phase differences at merger typically remain below a few portions of a radian, with the notable exception of the $(3,2)$ and $(4,2)$ modes. Notably, this fact remains true also for the $q=4$ simulation, which was run at a lower resolution with respect to the other systems.
Figure 19: Time domain comparison between the GR-Athena++ (red dashed) and SXS (black) waveforms for the $(\ell,m)=\{(2,2),(3,2),(4,4),(4,2)\}$ modes and $q=1$. Top panels show the real part of the waveform multipoles, while the bottom panels display the phase difference $\Delta\phi=\phi_{\ell m}^{\rm SXS}-\phi_{\ell m}^{\rm GRA}$ and the amplitude relative difference $\Delta A/A=1-A_{\ell m}^{\rm GRA}/A_{\ell m}^{\rm SXS}$. The vertical dashed-doted lines indicate the interval used for the alignment, while the vertical dashed lines mark the merger time. We find that while the $(2,2)$ and $(4,4)$ modes exhibit excellent agreement between the two waveforms, the $(3,2)$ and $(4,2)$ modes show larger discrepancies.
Figure 20: Time domain comparison between the GR-Athena++ (red dashed) and SXS (black) waveforms for the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(4,2),(5,5)\}$ modes and $q=3$. Top panels show the real part of the waveform multipoles, while the bottom panels display the phase difference $\Delta\phi=\phi_{\ell m}^{\rm SXS}-\phi_{\ell m}^{\rm GRA}$ and the amplitude relative difference $\Delta A/A=1-A_{\ell m}^{\rm GRA}/A_{\ell m}^{\rm SXS}$. The vertical dashed-doted lines indicate the interval used for the alignment, while the vertical dashed lines mark the merger time. Similarly to the $q=1$ case, modes with $\ell=m$ – including the $(5,5)$ mode – are in excellent agreement, while the $(3,2)$ and $(4,2)$ modes show larger discrepancies.
Figure 21: Time domain comparison between the GR-Athena++ (red dashed) and SXS (black) waveforms for the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(4,2),(5,5)\}$ modes and $q=4$. Top panels show the real part of the waveform multipoles, while the bottom panels display the phase difference $\Delta\phi=\phi_{\ell m}^{\rm SXS}-\phi_{\ell m}^{\rm GRA}$ and the amplitude relative difference $\Delta A/A=1-A_{\ell m}^{\rm GRA}/A_{\ell m}^{\rm SXS}$. The vertical dashed-doted lines indicate the interval used for the alignment, while the vertical dashed lines mark the merger time. Despite the lower resolution of the GRA waveforms with respect to other mass ratios (320 vs 384), the visual agreement and phase differences are comparable to those found with more refinement.

### v3

@@ Appendix F Time domain comparisons with SXS
In this appendix we report time domain comparisons similar to the one presented in the main text (Fig. 8) for all mass ratios considered in this work. The results are shown in Figs. 20, 21 and 22. Quantitative figures such as amplitude and phase differences are comparable with the ones discussed in the main portion of the text for the $q=2$ system. Phase differences at merger typically remain below a few portions of a radian, with the notable exception of the $(3,2)$ modes. Notably, this fact remains true also for the $q=4$ simulation, which was run at a lower resolution with respect to the other systems.
Figure 20: Time domain comparison between the GR-Athena++ (red dashed) and SXS (black) waveforms for the $(\ell,m)=\{(2,2),(3,2),(4,4)\}$ modes and $q=1$. Top panels show the real part of the waveform multipoles, while the bottom panels display the phase difference $\Delta\phi=\phi_{\ell m}^{\rm SXS}-\phi_{\ell m}^{\rm GRA}$ and the amplitude relative difference $\Delta A/A=1-A_{\ell m}^{\rm GRA}/A_{\ell m}^{\rm SXS}$. The vertical dashed-doted lines indicate the interval used for the alignment, while the vertical dashed lines mark the merger time. We find that while the $(2,2)$ and $(4,4)$ modes exhibit excellent agreement between the two waveforms, the $(3,2)$ mode shows larger discrepancies.
Figure 21: Time domain comparison between the GR-Athena++ (red dashed) and SXS (black) waveforms for the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes and $q=3$. Top panels show the real part of the waveform multipoles, while the bottom panels display the phase difference $\Delta\phi=\phi_{\ell m}^{\rm SXS}-\phi_{\ell m}^{\rm GRA}$ and the amplitude relative difference $\Delta A/A=1-A_{\ell m}^{\rm GRA}/A_{\ell m}^{\rm SXS}$. The vertical dashed-doted lines indicate the interval used for the alignment, while the vertical dashed lines mark the merger time. Similarly to the $q=1$ case, modes with $\ell=m$ – including the $(5,5)$ mode – are in excellent agreement, while the $(3,2)$ mode shows larger discrepancies.
Figure 22: Time domain comparison between the GR-Athena++ (red dashed) and SXS (black) waveforms for the $(\ell,m)=\{(2,2),(2,1),(3,3),(3,2),(4,4),(5,5)\}$ modes and $q=4$. Top panels show the real part of the waveform multipoles, while the bottom panels display the phase difference $\Delta\phi=\phi_{\ell m}^{\rm SXS}-\phi_{\ell m}^{\rm GRA}$ and the amplitude relative difference $\Delta A/A=1-A_{\ell m}^{\rm GRA}/A_{\ell m}^{\rm SXS}$. The vertical dashed-doted lines indicate the interval used for the alignment, while the vertical dashed lines mark the merger time. Despite the lower resolution of the GRA waveforms with respect to other mass ratios (320 vs 384), the visual agreement and phase differences are comparable to those found with more refinement.